OpenStudy (anonymous):

Solve each inequality. 6(c-1) < -18 -6/7y - 6 >= 42 3x +5 <= 2x - 8 42w > 2(w + 7)

OpenStudy (shamil98):

6(c-1) < -18 first start off by dividing both sides by 6. c - 1 < -3 add 1 to both sides. c < -2

OpenStudy (isaiah.feynman):

-6/7y-6>=42 Add 6 to both sides -6/7y>=48 Divide both sides by -6/7 y<=-56

OpenStudy (anonymous):

lol